OpenStudy (anonymous):

Find f (g(x)) if f(x) = 2x - 1 and g(x) = x + 2.

OpenStudy (anonymous):

f(x+2) = 2(x+2)-1 f(x+2) = 2x+1

OpenStudy (anonymous):

That should be 2x+3...

OpenStudy (anonymous):

so what is it ?

OpenStudy (anonymous):

2x+3. You just plug in the g(x) into the f(x). I made an arithmetic error in the first one

OpenStudy (anonymous):

thank you !